OpenStudy (ny,ny):

Lou wants to make a coffee mixture to sell. He is going to mix Sumatra coffee which costs $2.50 per pound with Columbian coffee which costs $3.75 per pound. He wants to make 50 pounds of mix and he wants the cost of the mix to be $3.35 per pound. How many pound of each will he need?

OpenStudy (ny,ny):

here's what my teacher gave as a study guide if it helps any better... define: x = ___ % y= ___% system: x+y= (total weight) %x+%y= %(total weight) **percent as decimals

hero (hero):

Okay, and why haven't you done it yet?

OpenStudy (ny,ny):

because what is x? and what is y?

OpenStudy (ny,ny):

total weight is 50, that i know. but is x 2.50 and is y 3.75? and where do i put 3.35 in the equation? as the total weight for the second equation?

hero (hero):

x + y = 50 2.50x + 3.75y = 3.35(50) Solve the system

OpenStudy (ny,ny):

use elimination?

hero (hero):

x + y = 50 2.50x + 3.75y = 167.50 Multiply first equation by 2.50: 2.50x + 2.50y = 125.00 2.50x + 3.75y = 167.50 Multiply the 1st equation by -1: -2.50x - 2.50y = -125.00 2.50x + 3.75y = 167.50 1.25y = 42.00 y = 42.00/1.25 y = 33.60 x = 16.40

hero (hero):

x = Samutra y = Columbian

OpenStudy (ny,ny):

wow that is very helpful... do i always have to put the total weight in the first equation multiplied by the total cost in the second equation? like when you did 3.35(50).

hero (hero):

Depends on the problem

OpenStudy (ny,ny):

mkay...

hero (hero):

I still have yet to meet someone who says mkay in real life. So far, it's all online.

OpenStudy (ny,ny):

yeah i guess :P

OpenStudy (ny,ny):

thanks you fer the help.

hero (hero):

You're welcome mkay

OpenStudy (ny,ny):

lol